Circular No. 6352

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996O IN MCG 03-41-115 B. Schmidt, Mount Stromlo and Siding Spring Observatories, e-mails:

“R. O. Evans reports his discovery of a supernova of mag ∼ 18 on a U.K. Schmidt R film taken by P. Cass on Mar. 21 as part of the U.K. Schmidt Supernova Search. SN 1996O is located at α = 16^{h}06^{m}00^{s}, δ = +18^{o}11′.8 (equinox 2000.0), 20′′ west and 2′′ south of the nucleus of MCG 03-41-115. The object was not visible on a similar film taken in Aug. 1995. A spectrum (range 550–900 nm) obtained with the Anglo-Australian Telescope on Mar. 23.7 by L. Germany (MSSO), B. Schmidt (MSSO) and R. Stathakis (Anglo-Australian Observatory), shows the object to be a type Ia supernova, ∼ 50 days past maximum. The redshift of the supernova, determined from narrow Hα emission in the host galaxy, is z = 0.03.”

COMET C/1995 Y1 (HYAKUTAKE) Ephemeris continuation (cf. IAUC 6293) from the orbital elements on

MPC 26723:

1996 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Mar. 18 20^{h}43.^{m}14 +21^{o}07.^{′}1 1.436 1.120 51.^{o}0 43.^{o}7  9.3
23 21 03.10 +23 20.3 1.492 1.151 50.4 41.9  9.5
28 21 21.88 +25 16.7 1.549 1.186 49.9 40.1  9.7
Apr. 2 21 39.48 +26 58.1 1.607 1.225 49.6 38.4  9.9
7 21 55.93 +28 26.6 1.665 1.268 49.4 36.9 10.1
12 22 11.27 +29 44.2 1.722 1.313 49.4 35.4 10.4
17 22 25.56 +30 52.5 1.777 1.362 49.6 34.2 10.6
22 22 38.84 +31 53.0 1.828 1.412 50.1 33.1 10.8
27 22 51.15 +32 46.8 1.877 1.464 50.7 32.1 11.0
May 2 23 02.55 +33 35.0 1.922 1.518 51.6 31.4 11.2
7 23 13.07 +34 18.3 1.963 1.574 52.8 30.7 11.4
12 23 22.77 +34 57.4 2.001 1.630 54.2 30.2 11.6
17 23 31.65 +35 32.9 2.034 1.687 55.8 29.7 11.8
22 23 39.74 +36 05.1 2.062 1.744 57.6 29.4 12.0
27 23 47.05 +36 34.2 2.087 1.802 59.7 29.0 12.2

Total visual magnitude estimates: Mar. 5.16 UT, 8.2 (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector); 9.16, 8.8 (Hornoch); 14.51, 8.7 (A. Hale, Cloudcroft, NM, 0.41-m reflector); 16.09, 8.5 (Hornoch); 20.11, 8.6 (Hornoch); 21.21, 8.7 (A. Pereira, Cabo da Roca, Portugal, 0.15-m reflector); 23.48, 9.1 (Hale).

1996 March 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Brian G. Marsden